OpenStudy (anonymous):

Let u = <-6, 3>, v = <1, 9>. Find u - v. <-15, 2> <-9,-8> <-5, 12> <-7, -6>

OpenStudy (praxer):

it is D, <-6-1,3-9> => <-7,-6>

OpenStudy (anonymous):

^ this + cute kitten picture!

OpenStudy (anonymous):

an you help me with this one as well

OpenStudy (anonymous):

Solve the triangle. A = 46°, a = 31, b = 27

OpenStudy (praxer):

Use the Law of Sines to determine B: sin(A)/a = sin(B)/b b sin(A)/a = sin(B) 27 sin(46°)/31 = sin(B) 0.625652176 ≈ sin(B) 38.8° ≈ B C = 180 - A - B ≈ 95.2° Use the Law of Sines again to determine C. You will find that c ≈ 42.9.

OpenStudy (anonymous):